# Pretraining-finetuning Framework for Efficient Co-design: A Case Study on Quadruped Robot Parkour

### v1

@@ Abstract
In nature, animals with exceptional locomotion abilities, such as cougars, often possess asymmetric fore and hind legs, with their powerful hind legs acting as reservoirs of energy for leaps. This observation inspired us: could optimize the leg length of quadruped robots endow them with similar locomotive capabilities? In this paper, we propose an approach that co-optimizes the mechanical structure and control policy to boost the locomotive prowess of quadruped robots. Specifically, we introduce a novel pretraining-finetuning framework, which not only guarantees optimal control strategies for each mechanical candidate but also ensures time efficiency. Additionally, we have devised an innovative training method for our pretraining network, integrating spatial domain randomization with regularization methods, markedly improving the network’s generalizability. Our experimental results indicate that the proposed pretraining-finetuning framework significantly enhances the overall co-design performance with less time consumption. Moreover, the co-design strategy substantially exceeds the conventional method of independently optimizing control strategies, further improving the robot’s locomotive performance and providing an innovative approach to enhancing the extreme parkour capabilities of quadruped robots.
@@ I INTRODUCTION
Quadruped animals exhibit exceptional mobility capabilities. For instance, a cougar can leap up to 5.5 meters high, and with a run-up, it can jump 12 meters. Similarly, a red kangaroo, about 1.8 meters tall, can easily leap forward more than 9 meters, quintupling its height. Imagine if legged robots could harness similar locomotive prowess, this could significantly boost their functionality, especially in challenging environments. It is noteworthy that quadruped mammals typically exhibit an asymmetric leg structure, a result of natural evolution that favors agile movement in their natural habitats [1]. This realization spurred our curiosity: Could adjusting the leg length of standard quadruped robots potentially unlock greater locomotive capabilities?
@@
Figure 1: Comparison of co-design methods. (a) Designing distinct control strategies tailored to various morphologies, which demands considerable computational resources. (b) Utilizing a generalized strategy as a substitute, which saves time but cannot guarantee optimality. (c) The algorithm we introduce starts by training a generalized model across varying leg lengths, followed by fine-tuning for specific morphology. The fine-tuning process requires merely a small percentage of steps typically necessary, striking a balance between performance and time-saving efficacy.
@@
The joint optimization of a robot’s mechanical structure and control policy is referred to as co-design, and currently, there are numerous works focused on the co-design of quadruped robots [2, 3, 4, 5, 6]. These studies are primarily confined to walking tasks. For instance, [2] explored the ideal proportion between upper and lower limbs, optimizing for objectives such as speed tracking and energy efficacy across varied terrains. [5] accentuated the generalizability of the transformer-based control strategy, asserting its adaptability across a myriad of leg length ratios. It also showcased that using a fixed control policy, solely morphological updates could enable a robot to step over a 10 $\mathrm{c}\mathrm{m}$ platform. [3, 4] improved the mobility of small-scale quadruped robots with limited degrees of freedom on flat ground through co-design approaches. However, the potential of robot co-design has not yet been fully realized in simple walking tasks. In more challenging control tasks like parkour tasks [7, 8, 9, 10], such as high jumping and long jumping, the significance of co-design becomes even more prominent. The physical capabilities of a robot and the scope of tasks it can execute are constrained by its hardware components. When the mechanical structure is not aptly designed, even the most ingeniously crafted motion control policies might struggle to fulfill the intended tasks. It is through a reciprocal enhancement of mechanical structure and control policies that the functional ceiling of a robot can be significantly elevated.
@@
The co-design of robots can be primarily divided into methods based on dynamic models [11, 12, 13] and methods based on deep reinforcement learning (DRL) [2, 3, 4, 5, 6, 14, 15, 1, 16, 17]. However, dynamic model based methods require a wealth of expert knowledge and engineering experience to construct dynamical models and design equations and inequality constraints. For robots with different mechanical structures, system identification [18] must also be incorporated for motion control, which can be exceedingly tedious and complicated. DRL-based methods typically frame the problem as a bi-level optimization issue, with the lower-level optimizing the control strategy for a specified morphology and the upper-level searching for the best morphology given the optimal control policy. Depending on how the control strategy is obtained, there are mainly two categories. The first involves training specific controllers for given candidate structures during the upper-level optimization, as shown in Fig. 1(a). For instance, [1] uses a tournament algorithm at the upper-level, employing the fitness of each candidate morphology as input to derive the optimal morphology. On the lower-level, it utilizes 1152 CPUs with DRL algorithms to learn the optimal control policies for various candidate morphologies, which is highly time-consuming and requires substantial computational power. In contrast, the second category focuses on developing a versatile, generalizable model [3, 2, 5, 6], capable of offering alternative strategies for various morphologies, as depicted in Fig. 1(b). For example, [3] employs a concurrent network architecture, consisting of both individual and population networks. At the lower-level, individual networks are trained for control policies of robots with specific morphology, while at the upper-level, the population network integrates all the training data from different morphologies to achieve a generalized strategy, hence saving time by avoiding the need for separate control policies for each candidate morphology. Similarly, [6] uses the teacher-student approach, and [5] leverages a transformer architecture, both training control strategies that can generalize across different morphologies, significantly time-efficient compared to the first category methods. However, this approach of substituting optimal strategies with generalized ones for different structures, while time-saving, cannot ensure optimality in specific morphologies, potentially impairing the efficacy of upper-level optimization decisions.
@@
In order to maintain optimality while also conserving computational power, in this paper, we introduced a pretraining-finetuning framework for quadruped robots’ co-design, with the goal of improving their parkour performance, as shown in Fig. 1(c). We first trained a parkour strategy for robots that is adaptable across a variety of leg lengths. Specifically, we integrated spatial domain randomization and regularization techniques. The former enables simultaneous training of thousands of robots with varying leg lengths, while the latter helps in reducing memory biases in the value network. The combination of them significantly enhances the generalizability of the pre-trained model. In the subsequent phase of morphology optimization, we fine-tuned this generalized model to suit each specific morphology candidate. Our experimental results indicate that merely 400 steps of fine-tuning (6.67% of the typical training steps) were sufficient for the new model to converge, delivering performance comparable to models specifically designed for each structure. In summary, the contributions of this paper are as follows:
@@ II Methodology
Figure 2: (a) The URDF model of the quadruped robot, where the solid rectangle represents the robot’s link, and the dashed ellipse represents the robot’s joint. The black part denotes fixed parameters, while the blue part denotes parameters that need to be changed. (b) The quadruped robot with different scaling parameters. For better comprehension, in this figure, the robots positioned further away have longer legs.
Figure 3: Overview of the proposed co-design pipeline. We initially pre-train a control strategy capable of generalizing across various morphologies. Subsequently, we embed the fine-tuning process into the morphological optimization phase, fine-tuning the pre-trained strategy for each set of candidate morphologies. This approach provides more accurate fitness values for morphological optimization. We employed the Proximal Policy Optimization (PPO) algorithm to update our policy, utilizing an asymmetric actor-critic architecture. Additionally, during the pre-training phase, we employed the technique of curriculum learning.
@@ II-A System Overview
The overall framework of our method is depicted in Fig. 3, including two phases: Pre-training (Section II-B) and Bayesian optimization (BO) with embedded fine-tuning (Section II-C). In the first phase, we develop a parkour policy that generalizes to varying leg lengths, serving as an initial model. During the second phase, the algorithm iteratively finds the optimal morphology using the fitness of candidate morphologies as prior knowledge. Since the pre-trained model might not be optimal for every specific morphology, we fine-tune it for better accuracy. This fine-tuning allows for more precise fitness calculations in BO, improving the co-design process’s overall effectiveness.
We frame the parkour control problem as a Markov Decision Process (MDP), which is defined by a tuple $\left\langle{\mathcal{S},\mathcal{A},P,r,\gamma}\right\rangle$, in which $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s_{t+1}|s_{t},a_{t})$ is the transition probability, $r(s_{t},a_{t})$ is the reward function, and $\gamma$ is the discount factor. At each timestep, the agent receives a state $s_{t}\in\mathcal{S}$ from the environment and selects an action $a_{t}\in\mathcal{A}$, guided by policy $\pi_{\theta}(a_{t}|s_{t})$. This action yields a reward $r_{t}$. The objective of DRL is to derive an optimal policy $\pi^{*}$ that maximizes the expected sum of discounted rewards ${\mathbb{E}}\left[{\sum_{t=0}^{T}{{\gamma^{t}}{r_{t}}}}\right]$.
Design of State and Action Space. The state comprises four parts: the proprioceptive state $x_{t}$, the explicit privileged state ${e^{\prime}_{t}}$, the privileged state $e_{t}$, the exteroceptive state $m_{t}$, and the historical proprioceptive state $h_{t}$. The $x_{t}$ includes the body’s angular velocity, roll angle, pitch angle, the difference between the body’s yaw angle and the target yaw angle, positions and velocities of joints, the actions in the previous step, and foot contact information. The $e^{\prime}_{t}$ is the body’s linear velocity. $e_{t}$ encompasses the robot’s mass, center of mass position, robot morphology parameters, friction factor, and motor strength parameters. $m_{t}$ consists of height sampling points. The action $a_{t}$ is the target angles for the robot’s twelve joints, which are converted into torques through a PD controller and then transmitted to the robot.
Design of Reward Function. Unlike previous works [19, 20, 21, 22], which utilized velocity tracking as the primary reward function item, manually specifying the robot’s linear velocity and its yaw angle in parkour tasks constraints the solution space of actions, leading to failure in performing parkour tasks. We designed our reward functions by referring to [9]. The primary reward function Goal Tracking, encourages the robot to move toward waypoint markers. When encountering an elevated platform, this reward function item motivates the robot to jump onto it rather than bypass it, as in obstacle avoidance tasks. The design of the reward function item is shown in Tab. VI of the Appendix.
@@ II-B1 Spatial Domain Randomization
Domain randomization emerges as a pivotal technique for augmenting the generalization capacity of DRL models. By sampling dynamic parameters within a specified reasonable range, this method significantly enhances the model’s robustness. This approach is primarily utilized to bridge the sim2real gap [23] in DRL. Unlike previous approaches [24, 21], our approach, utilizing domain randomization, is designed to obtain a pre-trained model with exceptional generalization abilities, thus establishing a robust foundation for subsequent fine-tuning phases. In traditional practices, domain randomization has involved the alteration of variables such as friction coefficients and motor strengths during training iterations, a process we refer to as Temporal Domain Randomization. However, altering the robot’s mechanical parameters, which are predetermined at the onset of the simulator, presents a significant challenge for modifications during training. To address this, we introduce the concept of Spatial Domain Randomization. Leveraging the extensive parallel capabilities of Isaac Gym [25], we implement this innovative approach by simultaneously training thousands of robots with varied leg lengths in a single environment. This approach allows us to develop a parkour control strategy with good generalization in just a few hours.
Similar to our previous work [4], we create robots with varying morphological parameters by modifying the unified robot description format (URDF) files of the quadruped robots. The URDF of a quadruped robot is illustrated in Fig. 2(a). In this paper, we primarily focus on altering the lengths of the robot’s shanks and thighs, as indicated by the blue parts in Fig. 2(a). For each link and joint requiring modification, the specific parameters to be changed are detailed in Tab. I. To ensure the robot’s stability, we make the corresponding parts of the robot’s left and right legs symmetrical. Therefore, the adjustable parameters are the scaling factors for the front thighs, front shanks, hind thighs, and hind shanks. When training the generalizable model, we allow the scaling factors to randomly sample within a feasible range, ${\xi_{i}}\sim\mathcal{U}(c_{min},c_{max})$. The sampled parameters ${\xi_{i}}$ are then used to construct robots with varied leg lengths, as illustrated in Fig. 2(b).
@@ II-B2 Regularizations
In the previous section, we introduced the spatial domain randomization method, which enables the algorithm to train using data from robots with varying leg lengths. As introduced in [26, 27], a value network trained across multiple agents is more likely to memorize the training data, leading to poor generalization to unvisited states. This not only hinders training performance but also negatively impacts testing performance in new environments. To mitigate this issue, we have implemented two types of regularization methods. The first method is Activation Regularization, which entails the inclusion of a regularization term in the critic’s loss function of PPO [28], as shown below:
[…]
where $\hat{A}_{t}^{\gamma}={\sum_{l=0}^{\infty}{(\gamma\lambda)}^{l}}\delta_{t+l}^{\gamma}$, this term provides an estimate of the advantage function at time step $t$ using a discount factor $\gamma$. $\lambda$ is the hyperparameter associated with Generalized Advantage Estimation (GAE) [29]. The temporal difference error $\delta_{t}^{\gamma}$ is defined as $r_{t}+\gamma V_{\phi}(s_{t+1})-V_{\phi}(s_{t})$. $\beta$ is a hyperparameter, which aims to penalize large value estimates, thereby promoting more consistent value estimation across different state-action pairs. Such a strategy potentially fosters generalization by reducing the impact of spurious approximation errors.
[…]
$J_{V}^{\rm{D}}(\phi)={\mathbb{E}_{{s_{t}}\sim\tau}}\left[{\frac{1}{2}{{\left({{V_{\phi}}({s_{t}})-\hat{R}_{t}^{{\gamma_{\rm{reg}}}}}\right)}^{2}}}\right]$ (2)
A lower discount factor reduces reliance on future rewards, making the algorithm focus on short-term gains. This approach may improve generalization by reducing variance.

### v2

@@ Abstract
In nature, animals with exceptional locomotion abilities, such as cougars, often possess asymmetric fore and hind legs. This observation inspired us: could optimizing the leg length of quadruped robots endow them with similar locomotive capabilities? In this paper, we propose an approach that co-optimizes the mechanical structure and control policy to boost the locomotive prowess of quadruped robots. Specifically, we introduce a novel pretraining-finetuning framework, which not only guarantees optimal control strategies for each mechanical candidate but also ensures time efficiency. Additionally, we have devised an innovative training method for our pretraining network, integrating spatial domain randomization with regularization methods, markedly improving the network’s generalizability. Our experimental results indicate that the proposed pretraining-finetuning framework significantly enhances the overall co-design performance with less time consumption. Moreover, the co-design strategy substantially exceeds the conventional method of independently optimizing control strategies, further improving the robot’s locomotive performance and providing an innovative approach to enhancing the extreme parkour capabilities of quadruped robots.
@@ I INTRODUCTION
Quadruped animals exhibit exceptional mobility capabilities. For instance, a red kangaroo, about 1.8 meters tall, can easily leap forward more than 9 meters, quintupling its height. It is noteworthy that quadruped mammals typically exhibit an asymmetric leg structure, a result of natural evolution that favors agile movement in their natural habitats [1]. This realization spurred our curiosity: Could adjust the leg length of standard quadruped robots potentially unlock greater locomotive capabilities?
@@
Figure 1: Comparison of co-design methods. (a) Designing distinct control strategies tailored to various morphologies. (b) Utilizing a generalized strategy as a substitute. (c) The algorithm we introduce starts by training a generalized model across varying leg lengths, followed by fine-tuning for specific morphology.
@@
The joint optimization of a robot’s mechanical structure and control policy is referred to as co-design, and currently, there are numerous works focused on the co-design of quadruped robots [2, 3, 4, 5, 6]. These studies are primarily confined to walking tasks. For instance, [2] explored the ideal proportion between thighs and shanks, optimizing for objectives such as speed tracking and energy efficacy across varied terrains. [5] accentuated the generalizability of the transformer-based control strategy, asserting its adaptability across a myriad of leg length ratios. It also showcased that using a fixed control policy, solely morphological updates could enable a robot to step over a 10 $\mathrm{c}\mathrm{m}$ platform. [3, 4] improved the mobility of small-scale quadruped robots with limited degrees of freedom on flat ground through co-design approaches. However, the potential of robot co-design has not yet been fully realized in simple walking tasks. In more challenging control tasks like parkour tasks [7, 8, 9, 10], such as high jumping and long jumping, the significance of co-design becomes even more prominent. The physical capabilities of a robot and the scope of tasks it can execute are constrained by its hardware components. When the mechanical structure is not aptly designed, even the most ingeniously crafted motion control policies might struggle to fulfill the intended tasks. It is through a reciprocal enhancement of mechanical structure and control policies that the functional ceiling of a robot can be significantly elevated.
@@
The co-design of robots can be primarily divided into methods based on dynamic models [11, 12, 13] and methods based on deep reinforcement learning (DRL) [2, 3, 4, 5, 6, 14, 15, 1, 16, 17]. However, dynamic model-based methods require a wealth of expert knowledge and engineering experience to construct dynamical models and design equations and inequality constraints. For robots with different mechanical structures, system identification [18] must also be incorporated for motion control, which can be exceedingly tedious and complicated. DRL-based methods typically frame the problem as a bi-level optimization issue, with the lower-level optimizing the control strategy for a specified morphology and the upper-level searching for the best morphology given the optimal control policy. Depending on how the control strategy is obtained, there are mainly two categories. The first involves training specific controllers for given candidate structures during the upper-level optimization, as shown in Fig. 1(a). For instance, [1] uses a tournament algorithm at the upper-level, employing the fitness of each candidate morphology as input to derive the optimal morphology. On the lower-level, it utilizes 1152 CPUs with DRL algorithms to learn the optimal control policies for various candidate morphologies, which is highly time-consuming and requires substantial computational power. In contrast, the second category focuses on the development of a versatile and generalizable model [3, 2, 5, 6], which is capable of providing alternative strategies applicable to various morphologies, as illustrated in Fig. 1(b). Specifically, [3] employs a concurrent network architecture, [6] adopts a teacher-student architecture, and [5] leverages a transformer architecture. However, this approach of substituting optimal strategies with generalized ones, while time-saving, cannot ensure optimality in specific morphologies, potentially impairing the efficacy of upper-level optimization decisions.
Figure 2: (a) The URDF model of the quadruped robot, where the solid rectangles represent links, and the dashed ellipses represent joints. The black parts denote fixed parameters, while the blue parts denote parameters that need to be changed. (b) The quadruped robot with different scaling parameters.
@@
In order to maintain optimality while also conserving computational power, in this paper, we introduced a pretraining-finetuning framework for quadruped robots’ co-design, with the goal of improving their parkour performance, as shown in Fig. 1(c). We first trained a parkour strategy that is adaptable across a variety of morphologies. Specifically, we integrated spatial domain randomization and regularization techniques. The former enables simultaneous training of thousands of robots with varying morphologies, while the latter helps in reducing memory biases in the value network. The combination of them significantly enhances the generalizability of the pre-trained model. In the subsequent phase of morphology optimization, we fine-tuned this generalized model to suit each specific morphology candidate. Our experimental results indicate that merely 400 steps of fine-tuning (6.67% of the typical training steps) were sufficient for the new model to converge, delivering performance comparable to models specifically designed for each structure. In summary, the contributions of this paper are as follows:
@@ II Methodology
Figure 3: Overview of the proposed co-design pipeline. We initially pre-train a control strategy capable of generalizing across various morphologies. Subsequently, we embed the fine-tuning process into the morphological optimization phase, fine-tuning the pre-trained strategy for each set of candidate morphologies. This approach provides more accurate fitness values for morphological optimization. We employed the Proximal Policy Optimization (PPO) algorithm to update our policy, utilizing an asymmetric Actor-Critic architecture.
@@ II-A System Overview
The overall framework of our method is depicted in Fig. 3, including two phases: Pre-training (Section. II-B) and Bayesian optimization (BO) with embedded fine-tuning (Section. II-C). We frame the parkour control problem as a Markov Decision Process (MDP), which is defined by a tuple $\left\langle{\mathcal{S},\mathcal{A},P,r,\gamma}\right\rangle$, in which $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s_{t+1}|s_{t},a_{t})$ is the transition probability, $r(s_{t},a_{t})$ is the reward function, and $\gamma$ is the discount factor. At each timestep, the agent receives a state $s_{t}\in\mathcal{S}$ from the environment and selects an action $a_{t}\in\mathcal{A}$, guided by policy $\pi_{\theta}(a_{t}|s_{t})$. This action yields a reward $r_{t}$. The objective of DRL is to derive an optimal policy $\pi^{*}$ that maximizes the expected sum of discounted rewards ${\mathbb{E}}\left[{\sum_{t=0}^{T}{{\gamma^{t}}{r_{t}}}}\right]$.
Design of State and Action Space. The state comprises four parts: the proprioceptive state $x_{t}$, which includes the body’s angular velocity, roll angle, pitch angle, the difference between the body’s yaw angle and the target yaw angle, positions and velocities of joints, the actions in the previous step, and foot contact information; the explicit privileged state ${e^{\prime}_{t}}$, which refers to the body’s linear velocity; the privileged state $e_{t}$, which encompasses the robot’s mass, center of mass position, robot morphology parameters, friction factor, and motor strength parameters; the exteroceptive state $m_{t}$, which consists of height sampling points; and the historical proprioceptive state $h_{t}$. The action $a_{t}$ is the target angles for the robot’s twelve joints, which are converted into torques through a PD controller and then transmitted to the robot.
Design of Reward Function. Unlike previous works [19, 20, 21, 22], which utilized velocity tracking as the primary reward function item, manually specifying the robot’s linear velocity and its yaw angle in parkour tasks constraints the solution space of actions, leading to failure in performing parkour tasks. We designed our reward functions by referring to [9]. The primary reward function Goal Tracking, encourages the robot to move toward waypoint markers. When encountering an elevated platform, this reward function item motivates the robot to jump onto it rather than bypass it, as in obstacle avoidance tasks. The design of the reward function item is shown in Section. -A of the Appendix.
@@ II-B1 Spatial Domain Randomization
Domain randomization emerges as a pivotal technique for augmenting the generalization capacity of DRL models[23, 21]. By sampling dynamic parameters within a specified reasonable range, this method significantly enhances the model’s robustness. In traditional practices, domain randomization has involved altering variables such as friction coefficients and motor strengths during training to bridge the sim2real gap [24], a process we refer to as Temporal Domain Randomization. However, in our task, the structural parameters of the robot are determined when the simulator is loaded. Any changes to these parameters necessitate reloading the simulator. Making such modifications in every training epoch would undoubtedly significantly slow down the algorithm’s training speed. To address this, we introduce the concept of Spatial Domain Randomization. Leveraging the extensive parallel capabilities of Isaac Gym [25], we implement this innovative approach by simultaneously training thousands of robots with varied morphologies in a single environment. This approach allows us to develop a parkour control strategy with good generalization in just a few hours.
Similar to our previous work [4], we create robots with varying morphological parameters by modifying the unified robot description format (URDF) files of the quadruped robots. The URDF of a quadruped robot is illustrated in Fig. 2(a). Although our method is applicable to various structural parameters of the robot, we have chosen to focus solely on the thigh and shank lengths for a clearer analysis of the optimization results, as indicated by the blue parts in Fig. 2(a). For each link and joint requiring modification, the specific parameters to be changed are detailed in Tab. I. To ensure the robot’s stability, we make the corresponding parts of the robot’s left and right legs symmetrical. Therefore, the adjustable parameters are the scaling factors for the front thighs, front shanks, hind thighs, and hind shanks. When training the generalizable model, we allow the scaling factors to randomly sample within a feasible range, ${\xi_{i}}\sim\mathcal{U}(c_{min},c_{max})$. The sampled parameters ${\xi_{i}}$ are then used to construct robots with varied leg lengths, as illustrated in Fig. 2(b). It’s worth noting that when the length of a robot’s limbs changes, the corresponding PD parameters must also be adjusted. Following the approach outlined in [26], we utilize a polynomial to calculate the corrected proportional factors as follows:
${\eta_{i}}=a\times\xi_{i}^{3}+b\times\xi_{i}^{2}+c\times{\xi_{i}}+d$ (1)
The constants $a$, $b$, and $c$ are manually specified. Since we modified the thigh and shank, which correspond to the HipY and Knee joints shown in Fig. 2(a), we apply the correction factor ${\eta_{i}}$ to the PD parameters of these eight joints.
@@ II-B2 Regularizations
The proposed spatial domain randomization method enables the algorithm to train using data from robots with varying leg lengths. However, as introduced in [27, 28], a value network trained across multiple agents is more likely to memorize the training data, leading to poor generalization to unvisited states. This not only hinders training performance but also negatively impacts testing performance in new environments. To mitigate this issue, we employed a Discount Regularization approach, which adjusts the discount factor to a smaller value, denoted as $\gamma_{\text{reg}}$, where $10<\gamma_{\text{reg}}<\gamma<$ 1. The modification of the discount factor influences the calculation of the advantage function and temporal difference in the PPO algorithm, thereby affecting the computation of the Actor and Critic loss functions. The new formulations for the advantage function and temporal difference are as follows:
$A_{t}^{{\gamma_{{\rm{reg}}}}}=\Sigma_{l=0}^{\infty}{\left({{\gamma_{{\rm{reg}}}}\lambda}\right)^{l}}\delta_{t+l}^{{\gamma_{{\rm{reg}}}}}$ (2)
$\delta_{t}^{{\gamma_{{\rm{reg}}}}}={r_{t}}+{\gamma_{{\rm{reg}}}}{V_{\phi}}\left({{s_{t+1}}}\right)-{V_{\phi}}\left({{s_{t}}}\right)$ (3)
where $\lambda$ is the hyperparameter used in Generalized Advantage Estimation (GAE)[29]. Theoretically, a lower discount factor decreases the reliance on future rewards, allowing the algorithm to concentrate more on short-term gains. This approach can enhance generalization by reducing variance. For the mathematical proof of this method, please refer to the Section. -B of the Appendix.